A motor can running at the rate of 7ms inverse is stopped by the breaks in 15m.then the ratio between the frictional force and weight of the car is

Grade:12th pass

1 Answers

Saurabh Koranglekar
3 years ago
Dear student

F = m*a

v^2 = u^2 +2aS

a= 49/30 m per sec^2

Regards